OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

the slope of your line is \[\frac{4}{3}\] so that is the slope of a parallel line. now use \[y=y_1=m(x-x_1)\] with \[x_1=-3,y_1=4,m=\frac{4}{3}\]

OpenStudy (anonymous):

First thing to do is write the line in slope intercept form so you know what the slope is.\[y = \frac{4}{3}x + 4\] In order for a line to be parallel to another they must have the same slope. So the slope of the line parallel to the line given is m = 4/3. Now we use the point-slope formula to write the equation. \[y - 4 = \frac{4}{3}(x + 3)\] Solve for y and you get \[y = \frac{4}{3}x + 8\]